OpenStudy (anonymous):

divide 7/8 and 7/40 simplify your answer

OpenStudy (anonymous):

do you mean divide 7/8 by 7/40?

OpenStudy (anonymous):

just multiply 7/8 by 40/7

OpenStudy (anonymous):

280/56=5

OpenStudy (anonymous):

yeah divide 7/8 n 7/40

OpenStudy (anonymous):

okay just flip the 7/40 to 40/7 and multiply

OpenStudy (anonymous):

40/7 is called the reciprocal

OpenStudy (anonymous):

im not sure how to do this doesnt the denomiantor have to be the same?

OpenStudy (anonymous):

when you divide two fractions the best way to do it is to multiply the first fraction by the reciprocal of the second fraction.

OpenStudy (anonymous):

ok and after that

OpenStudy (anonymous):

then just multiply both numerators and denomintors like I did and simplify

OpenStudy (anonymous):

7*40=280 8*7=56 so your answer is 280/56= 5

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

If this is what you mean, divide 7/8 by 7/40, then here it is: \( \dfrac{7}{8} \div \dfrac{7}{40} \) Keep the same first fraction and multiply it by the reciprocal of the second fraction (to get the reciprocal of a fraction, just flip it): \( =\dfrac{7}{8} \times \dfrac{40}{7} \) Now multiply the numerators together and the denominators together: \( \dfrac{7 \times 40}{8 \times 7}\) Now reduce before mutiplying the numbers: \( =\dfrac{7 \times 5 \times 8}{8 \times 7}\) \( =\dfrac{\cancel{7} \times 5 \times \cancel{8}}{\cancel{8} \times \cancel{7} 1 }\) \(=5\)

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm